OpenStudy (anonymous):

graph help f(x) and g(x) x=3 f(x)=2 g(x)=-4 g'(x)=1/5 h(x)=2g(x)+1/(x)-f(x)g(x)

OpenStudy (anonymous):

compute h'(3)=

OpenStudy (anonymous):

help please @wio

zepdrix (zepdrix):

\[\large h(x)=\frac{2g(x)+1}{x-f(x)g(x)}\]Is this what h(x) loos like?

OpenStudy (anonymous):

yes sure is

zepdrix (zepdrix):

So to find \(\large h'(x)\) we'll have to apply the quotient rule.

OpenStudy (anonymous):

ok, g(x)f(x)'-f(x)g(x)'/g(x)^2?

zepdrix (zepdrix):

It's probably not a good idea to think of the definition of the quotient rule in terms of \(\large f\) and \(\large g\) since that's going to confuse you in this problem :) lol Because in this problem our \(\large f\) is \(\large 2g(x)+1\). While our \(\large g\) is \(\large x-f(x)g(x)\). Understand? :x

OpenStudy (anonymous):

so far yes,

zepdrix (zepdrix):

\[\large h(x)=\frac{u}{v}\]\[\large h'(x)=\frac{\color{royalblue}{u'}v-u\color{royalblue}{v'}}{v^2}\] \[\large h'(x)=\frac{\color{royalblue}{\left[2g(x)+1\right]'}\left[x-f(x)g(x)\right]-\left[2g(x)+1\right]\color{royalblue}{\left[x-f(x)g(x)\right]'}}{\left[x-f(x)g(x)\right]^2}\] So here is our setup for quotient rule.

zepdrix (zepdrix):

It's a doozy. I tried to make it a little easier to read by changing the color of the ones we'll need to take a derivative of.

OpenStudy (anonymous):

its really easy to see when you colored it thank you.

OpenStudy (anonymous):

what is the next step?

zepdrix (zepdrix):

So all we've done so far is setup the quotient rule. We now need to take the derivative of each blue part. Let's try to do it piece by piece.

zepdrix (zepdrix):

So here is the upper left of our fraction.\[\large \color{royalblue}{\left[2g(x)+1\right]'}\left[x-f(x)g(x)\right]\]Taking the derivative gives us,\[\large \color{royalblue}{\left[2g'(x)+0\right]}\left[x-f(x)g(x)\right]\] Understand how I did that? We don't actually know what g(x) is, so we just throw a prime on it, to show that we've taken its derivative.

OpenStudy (anonymous):

Yes one is constant. so it became zero? please help me solve this question. Its my homework that is due tomorrow

OpenStudy (anonymous):

I want to get used to solve this question

zepdrix (zepdrix):

So the upper left of our fraction will simplify to,\[\large \left[2g'(x)\right]\left[x-f(x)g(x)\right]\] Now looking at the upper right of our fraction, this one is going to be a little tricky. We have to apply the product rule to the f and g part. \[\large -\left[2g(x)+1\right]\color{royalblue}{\left[x-f(x)g(x)\right]'}\] \[\large -\left[2g(x)+1\right]\color{royalblue}{\left[1-\color{orangered}{\left[f'(x)g(x)+f(x)g'(x)\right]}\right]}\]The x gave us 1, the product rule on f and g is in orange.

zepdrix (zepdrix):

The brackets around the orange are really important since there was a negative in front of the f(x)g(x), it will need to be distributed through.

OpenStudy (anonymous):

So far i understand. you explain very clearly

zepdrix (zepdrix):

So we've taken all the derivatives that we needed to take. This gives us an h'(x) of, \[\large \frac{\left[2g'(x)\right]\left[x-f(x)g(x)\right]-\left[2g(x)+1\right]\left[1-f'(x)g(x)-f(x)g'(x)\right]'}{\left[x-f(x)g(x)\right]^2}\]

zepdrix (zepdrix):

Quite the ugly looking problem.. wow your teacher is mean :d

zepdrix (zepdrix):

They want us to evaluate h'(x) at x=3. Which will look like this,\[\large \frac{\left[2g'(\color{#CC0033}{3})\right]\left[x-f(\color{#CC0033}{3})g(\color{#CC0033}{3})\right]-\left[2g(\color{#CC0033}{3})+1\right]\left[1-f'(\color{#CC0033}{3})g(\color{#CC0033}{3})-f(\color{#CC0033}{3})g'(\color{#CC0033}{3})\right]'}{\left[\color{#CC0033}{3}-f(\color{#CC0033}{3})g(\color{#CC0033}{3})\right]^2}\] And they gave us a bunch of info to fill in the blanks. \[\large f(3)=2\]\[\large g(3)=-4\]\[\large g'(3)=1/5\]

zepdrix (zepdrix):

But we have a small problem. In order to solve this, we need to know what \(\large f'(3)\) is. Did you forget to paste some of the info?

OpenStudy (anonymous):

haha really! um there is a graph like this: x f'(x) g(x) g'(x) 3 2 -2/3 1/5

OpenStudy (anonymous):

it says compute h'(3)=

zepdrix (zepdrix):

Oh so that was a small typo on your part I guess c: \(\large f'(3)=2\). But then we still need \(\large f(3)\) grrr -_- We can simplify this pretty far down, we won't be able to get a numerical answer though, no big deal.

zepdrix (zepdrix):

Oh I forgot to erase the prime on the upper right fraction, don't let that confuse you please :O That should be gone.

OpenStudy (anonymous):

hmm? which part?

zepdrix (zepdrix):

\[\large \frac{\left[2g'(x)\right]\left[x-f(x)g(x)\right]-\left[2g(x)+1\right]\left[1-f'(x)g(x)-f(x)g'(x)\right]\color{red}{'}}{\left[x-f(x)g(x)\right]^2}\] The one in red here. We have taken the derivative of everything inside. So the prime on the outside should disappear.

OpenStudy (anonymous):

so far im with you

zepdrix (zepdrix):

Grr it got cut off, imma have to paste it smaller.

zepdrix (zepdrix):

\[\frac{\left[2(1/5)\right]\left[3-f(3)(-2/3)\right]-\left[2(-2/3)+1\right]\left[1-(2)(-2/3)-f(3)(1/5)\right]}{\left[3-f(3)(-2/3)\right]^2}\]

OpenStudy (anonymous):

Thank you,

OpenStudy (anonymous):

i can see the whole thing now :)

zepdrix (zepdrix):

If you're confused about what I did, just look at the first set of brackets, maybe that will help you understand. \(\large \left[2g'(3)\right]\) became \(\large \left[2(1/5)\right]\) Because \(\large g'(3)=1/5\)

zepdrix (zepdrix):

See the problem we have? We have a big ugly mess since we weren't given information about f(3).

OpenStudy (anonymous):

yes I got it.:)

zepdrix (zepdrix):

Otherwise we would be able to simplify it down to just a number.

zepdrix (zepdrix):

Oh well :c heh

OpenStudy (anonymous):

Cool thanks for helping me.

OpenStudy (anonymous):

i have one question, may i ask?

zepdrix (zepdrix):

Yes, if it's not another 30 minute problem c: lolol

OpenStudy (anonymous):

yes, Thank you! its \[w=\frac{ \sqrt(t) }{ 2t^2+t3^t }\] find dw / dt

zepdrix (zepdrix):

\[\large w=\frac{\sqrt t}{2t^2+t3^t}\] Before taking this derivative, let's do some side work. Do you know the derivative of this term?\[\large 3^t\]

OpenStudy (anonymous):

something to do with In?

zepdrix (zepdrix):

lol, yes something to do with natural log :) We don't have time to go through the steps of that separately. So i'll just remind you of the rule. \[\large f(x)=e^x\]\[\large f'(x)=e^x\]Anytime the base is NOT e, you have to multiply by the natural log of the base.\[\large g(x)=2^x\]\[\large g'(x)=2^x(\ln 2)\] Just keep that in mind for when it comes up in this problem.

zepdrix (zepdrix):

\[\large w=\frac{\sqrt t}{2t^2+t3^t}\] \[\large w'=\frac{\color{royalblue}{(\sqrt t)'}\left[2t^2+t3^t\right]-\sqrt t\color{royalblue}{\left[2t^2+t3^t\right]'}}{\left[2t^2+t3^t\right]^2}\] Again, quotient rule. And all we really need to worry about is the blue parts.

zepdrix (zepdrix):

You lied! This is an evil problem -_- just like the last one. lolol

OpenStudy (anonymous):

I am so sorry...

OpenStudy (anonymous):

I cant even tell weather its not evil or not since I have no idea

OpenStudy (anonymous):

i mean not evil or evil

zepdrix (zepdrix):

If you don't have the derivative of sqrt memorized, then you'll want to change it to a rational expression so you can apply the power rule.

zepdrix (zepdrix):

Do you know that derivative? c: It's a good one to remember.

OpenStudy (anonymous):

power rule? n-1? and bring one number front

zepdrix (zepdrix):

yes... but do you understand how to take the derivative of the square root? :o

OpenStudy (anonymous):

No...

zepdrix (zepdrix):

\[\large \sqrt t \qquad = \qquad t^{1/2}\]You can take the derivative from this point by applying the power rule.

OpenStudy (anonymous):

oh i see. so it should 1/2t^(-1/2)

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

ok, could you write the whole thing?

OpenStudy (anonymous):

please dont go @zepdrix

zepdrix (zepdrix):

write the whole thing? :3 pshhhh ill bet you can do that XD Let's look at the other blue part a moment. We'll have to apply the product rule again.

OpenStudy (anonymous):

ok,i have been struggling with this problem for 1hr and half.

zepdrix (zepdrix):

\[\large \left[2t^2+t3^t\right]' \qquad = \qquad \left[4t+(t)'3^t+t(3^t)'\right]\]So here what i've done is setup the product rule, we'll still need to evaluate it.

OpenStudy (anonymous):

yes this is what i am struggling with!

zepdrix (zepdrix):

\[\large \left[4t+1\cdot3^t+t3^t(\ln 3)\right]\]Understand how that worked out?

OpenStudy (anonymous):

yes, so far. you used the rule that you explained before right

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

ok, whats next step

zepdrix (zepdrix):

So we were taking the derivative of each piece separately. Now just put the pieces back in. \[\large \frac{1/2t^{-1/2}\left[2t^2+t3^t\right]-\sqrt t\left[4t+1\cdot3^t+t3^t(\ln 3)\right]}{\left[2t^2+t3^t\right]^2}\] You could probably simplify it a bit.. but no big deal..

OpenStudy (anonymous):

how you simpligy Sqrt(t) and In3 part?

OpenStudy (anonymous):

i would appreciate it if you could show me how to simplify the whole thing

zepdrix (zepdrix):

Generally with these types of problems its better to not simplify them too far unless your teacher asks you to. They usually just want to see that you know how to take the derivative. It's not worth going any further really. :o

OpenStudy (anonymous):

Ok, thank you.

OpenStudy (anonymous):

you are the best math teacher :)

zepdrix (zepdrix):

lol haf good night c:

OpenStudy (anonymous):

thank you for your time night night!